Question 4: Use Chinese Remainder Theorem to find an integer a such that 4/ a +1,9/a +2, 25/a +3. 4
Answer #1

Given Alati 51442 25/a13 Note a = b (modny narb ar arbeo (modny > 941 zo moda azt (mo14) 1983 moda alat2 Ako 25 lat G+2E o moHere gud(4,9=ged (9.250 92d (25.91 Let m= 4x9125- guo Mi- М. go 45 4 225 M2 = M = @ m3= M. 36 Ya Now Find Y, Y Y,MI= 225 Y= IY,=1 Ч, = | 92 - 19 a= (Mi Yi Xi fм, ч, X, t Hз 3 X fмә4 0) 4- (3854 3 flvoyi x1 + 34 (Ч. 2.2) (Yo405) 4= (G) - — 20 е (1 509